# Kekulé order from diffuse nesting near higher-order Van Hove points

### v1

@@ B $K$-order stability analysis
Figure SM 6: Fermi surface nesting of the breathing configuration with nesting vector $\boldsymbol{q}=K$ at HOVHS filling with different broadening values $T$. When $T$ becomes infinitesimal the nesting condition vanishes.
@@
Upon lowering the FS broadening $T$, the $\Gamma$-peak in the particle-hole susceptibilty surpasses the apex at $\boldsymbol{q}=K$. Visualizing the FS overlap of a $K$-shifted FS for different values of $T$ (Fig. SM 6) in the breathing configuration elucidates this observation. In the absence of broadening, i.e. a weak-coupling scenario, the overlaps vanish because of non-matching curvatures of the FS regions.
@@
In the case of $K$-ordering, the appropriate symmetry group is no longer $C_{3v}$ of the original lattice, but the extended pointgroup $C_{3v}^{\prime\prime}$ of the enlarged $\sqrt{3}\times\sqrt{3}$ unit cell. When comparing the transformation behavior of the two states in Fig. SM 7 with the character table Table SI, the order parameter can be labeled by the $E_{3}^{\prime}$ irrep.
Figure SM 7: Degenerate leading FRG eigenvectors in the Kekulé phase at $K$ and ${K^{\prime}}$ up to nearest neighbor bond distances. The size (color) of the sites and bonds indicate the absolute value (complex phase) of the order parameter $\Delta^{K^{(\prime)}}$. The original (extended) Wigner-Seitz cell is indicated in grey (black).
For a symmetry group with multiple equivalent rotation centers like $C_{3v}$ (equivalent to the wallpaper group P3m1), the angular momentum for the center of mass coordinate of the order becomes dependent on the choice of the rotation center. E.g. if we choose the $C_{3}$ center in Fig. SM 7 as the center of a strong bond triangle, the obtained order acquires a trivial transformation behavior under rotation in contrast to the phase winding around the center of the hexagon.
Table SI: Character table of the extended pointgroup $C_{3v}^{\prime\prime}$ of the enlarged $\sqrt{3}\times\sqrt{3}$ unit cell of the breathing kagome lattice. $E$ is the identity operation, $t_{a_{1,2}}$ label translations with the primitive lattice vectors of the original kagome lattice, $C_{3}^{(2)}$ positive (negative) rotations by $32\pi/$ 3 and $\sigma_{1,2,3}$ are the three symmetry equivalent mirrors through the sides of each small triangle. We use the center of the hexagonal plaquettes as the symmetry center and note that a different rotation center leads to a permutation of the characters of the irreps $E_{i}^{\prime}$.
@@
$\bar{\Delta}^{K}_{o_{1}o_{2}}(k)=\Delta^{{K^{\prime}}}_{o_{2}o_{1}}(k+K)\ .$ (SM 4)
@@
From the FRG analysis, we know that the non-vanishing expectation value in the symmetry broken state is given by $\langle{\hat{c}^{\dagger}}_{k+q}{\hat{c}^{\vphantom{{\dagger}}}}_{k}\rangle$ with $q\in\{K,{K^{\prime}}\}$ and we can write down the corresponding mean field decoupling for the interacting Hamiltonian
$\begin{split}H=&\sum_{ko_{1}o_{2}}\mathcal{H}^{0}_{o_{1}o_{2}}(k){\hat{c}^{\dagger}}_{ko_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}+\frac{1}{N}\sum_{qk{k^{\prime}}\{o_{i}\}}V_{o_{1}o_{2}o_{3}o_{4}}(k,{k^{\prime}},q){\hat{c}^{\dagger}}_{k+qo_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}{\hat{c}^{\dagger}}_{{k^{\prime}}-qo_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{{k^{\prime}}o_{4}}\\ \simeq&\sum_{ko_{1}o_{2}}\mathcal{H}^{0}_{o_{1}o_{2}}(k){\hat{c}^{\dagger}}_{ko_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}+\frac{1}{N}\sum_{q\in\{K,{K^{\prime}}\}}\sum_{k{k^{\prime}}\{o_{i}\}}V_{o_{1}o_{2}o_{3}o_{4}}(k,{k^{\prime}},q)\Big{[}\langle{\hat{c}^{\dagger}}_{k+qo_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}\rangle{\hat{c}^{\dagger}}_{{k^{\prime}}-qo_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{{k^{\prime}}o_{4}}+{\hat{c}^{\dagger}}_{k+qo_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}\langle{\hat{c}^{\dagger}}_{{k^{\prime}}-qo_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{{k^{\prime}}o_{4}}\rangle\Big{]}+C\\ =&\sum_{ko_{1}o_{2}}\mathcal{H}^{0}_{o_{1}o_{2}}(k){\hat{c}^{\dagger}}_{ko_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}+2\sum_{q\in\{K,{K^{\prime}}\}}\sum_{k\{o_{i}\}}\Delta^{q}_{o_{1}o_{2}}(k){\hat{c}^{\dagger}}_{k+qo_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}+C=H_{\mathrm{MF}}\ ,\end{split}$ (SM 5)
@@
$\Delta^{q}_{o_{1}o_{2}}(k)=\frac{1}{N}\sum_{{k^{\prime}}o_{3}o_{4}}V_{o_{1}o_{2}o_{3}o_{4}}(k,{k^{\prime}},q)\langle{\hat{c}^{\dagger}}_{{k^{\prime}}-qo_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{{k^{\prime}}o_{4}}\rangle\ .$ (SM 6)
@@
$\begin{split}C=&-\frac{1}{N}\sum_{q\in\{K,{K^{\prime}}\}}\sum_{k{k^{\prime}}\{o_{i}\}}V_{o_{1}o_{2}o_{3}o_{4}}(k,{k^{\prime}},q)\langle{\hat{c}^{\dagger}}_{k+qo_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{ko_{2}}\rangle\langle{\hat{c}^{\dagger}}_{{k^{\prime}}-qo_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{{k^{\prime}}o_{4}}\rangle\\ =&-\frac{1}{N}\sum_{q\in\{K,{K^{\prime}}\}}\sum_{k{k^{\prime}}\{o_{i}\}}V_{o_{1}o_{2}o_{3}o_{4}}^{-1}(k,{k^{\prime}}q)\Delta_{o_{1}o_{2}}^{-q}(k)\Delta_{o_{3}o_{4}}^{q}({k^{\prime}})\end{split}$ (SM 7)
@@
Using the spinor notation ${\hat{\psi}^{\dagger}}_{ko_{1}}=\big{(}{\hat{c}^{\dagger}}_{ko_{1}},{\hat{c}^{\dagger}}_{k+Ko_{1}},{\hat{c}^{\dagger}}_{k+{K^{\prime}}o_{1}}\big{)}$, the MF Hamiltonian can be written as
$H_{\mathrm{MF}}=\frac{1}{3}\sum_{ko_{1}o_{2}}{\hat{\psi}^{\dagger}}_{ko_{1}}\begin{pmatrix}\mathcal{H}^{0}_{o_{1}o_{2}}(k)&2\Delta_{o_{1}o_{2}}^{{K^{\prime}}}(k+K)&2\Delta_{o_{1}o_{2}}^{K}(k+{K^{\prime}})\\ 2\Delta_{o_{1}o_{2}}^{K}(k)&\mathcal{H}^{0}_{o_{1}o_{2}}(k+K)&2\Delta_{o_{1}o_{2}}^{{K^{\prime}}}(k+{K^{\prime}})\\ 2\Delta_{o_{1}o_{2}}^{{K^{\prime}}}(k)&2\Delta_{o_{1}o_{2}}^{K}(k+K)&\mathcal{H}^{0}_{o_{1}o_{2}}(k+{K^{\prime}})\end{pmatrix}{\hat{\psi}^{\vphantom{{\dagger}}}}_{ko_{2}}+C=\frac{1}{3}\sum_{kn}\xi_{n}(k){\hat{\psi}^{\dagger}}_{kn}{\hat{\psi}^{\vphantom{{\dagger}}}}_{kn}+C$ (SM 8)
and we obtain the quasi-particle spectrum $\xi_{n}(k)$ within the ordered phase by diagonalizing the Hamiltonian. We note that the symmetry constraint of Eq. SM 4 guarantees a hermitian MF Hamiltonian. Exploiting the fermionic nature of the quasi-particles, i.e. ${\hat{\psi}^{\dagger}}_{kn}{\hat{\psi}^{\vphantom{{\dagger}}}}_{kn}=\{0,1\}$, the free energy can by directly calculated as
$F=-T\ln(Z)=-T\ln\Big{(}\prod_{kn}\sum_{n_{kn}=0}^{1}\mathrm{e}^{-\frac{\xi_{n}(k)}{3T}}\Big{)}+C=-T\sum_{kn}\ln\Big{(}1+\mathrm{e}^{-\frac{\xi_{n}(k)}{3T}}\Big{)}+C\ .$ (SM 9)
@@
$\Delta_{o_{1}o_{2}}(k)=\mathrm{e}^{\mathrm{i}\phi_{K}}A_{K}\Delta^{K}_{o_{1}o_{2}}(k)+\mathrm{e}^{\mathrm{i}\phi_{{K^{\prime}}}}A_{{K^{\prime}}}\Delta^{{K^{\prime}}}_{o_{1}o_{2}}(k)\ .$ (SM 10)
Since Eq. SM 4 is not enforced by TRS directly, but is rather a direct consequence of hermiticity and $K$ and ${K^{\prime}}$ being time reversal partners, this requirement persists in the symmetry broken phase. The order parameter in the $K$-phase features a finite site-local, and hence momentum independent, component. Thereby Eq. SM 4 directly implies $A_{K}=A_{{K^{\prime}}}$ and $\phi_{K}=-\phi_{{K^{\prime}}}$ and we need to optimize the free energy w.r.t. the simplified order parameter
$\Delta_{o_{1}o_{2}}(k)=A\big{[}\mathrm{e}^{\mathrm{i}\phi}\Delta^{K}_{o_{1}o_{2}}(k)+\mathrm{e}^{-\mathrm{i}\phi}\Delta^{{K^{\prime}}}_{o_{1}o_{2}}(k)\big{]}\ .$ (SM 11)
A MF analysis after an FRG flow technically requires a regularization of the divergent part of the two particle vertex via an inverse Bethe-Salpeter resummation [133, 134]. With this approach, all degrees of freedom are restored and the correct symmetry breaking on the mean field level is ascribed to a modified bare interaction. As apparent from Eq. SM 5, the effective two-particle interaction only enters the term $C$, which merely determines the absolute amplitude of the order parameter as a function of temperature, but cannot distinguish between different linear combinations within the degenerate manifold. Hence, we only need to evaluate the kinetic term in Eq. SM 9 as a function of $\phi$ for different values of $T\lesssim T_{\mathrm{C}}$ and $A$. This can be done for the original theory without any further restrictions [105].
The free energy in Fig. SM 8 shows three distinct minima at $\phi\in\{\pi/3,\pi,5\pi/3\}$ for all values of $A$ and $T$, which are related by a translation by an original primitive translation. Hence, all minima can be mapped back to the same state by a redefinition of the unit cell. The obtained state is depicted in Fig. 2 in the main text.
Figure SM 8: Phase dependence of the free energy for different inverse temperatures $\beta=1/T$ at fixed amplitude $A=t$ (solid lines). $F$ was calculated according to Eq. SM 9 on a $1000\times 1000$ momentum grid in the downfolded BZ. The dependence can be fitted by taking into account third and sixth order terms in the GL expansion in Eq. SM 14(dashed lines).
@@ A Complementary Ginzburg-Landau analysis
$F=\alpha\Delta^{K}\Delta^{K^{\prime}}+\gamma(\Delta^{K}\Delta^{K}\Delta^{K}+\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}})+\mathcal{O}(\Delta^{4})\ ,$ (SM 12)
where the summation over the internal momentum structure of $\Delta^{K^{(\prime)}}(k)$ has been carried out and incorporated in the GL coefficients $\alpha-\gamma$. While $\alpha\propto(T-T_{\mathrm{C}})$ determines the absolute size of the order parameter $A$ close to $T_{\mathrm{C}}$, the third order term $\propto\mathrm{Re}(\gamma)\cos(3\phi)-\mathrm{Im}(\gamma)\sin(3\phi)$ is responsible for fixing the overall phase of the condensate [135].
For $\beta=1/T=10$, i.e. small size of the order parameter, where the expansion in Eq. SM 12 is still valid, the phase dependence of the free energy can be appropriately fitted by setting $\mathrm{Re}(\gamma)>0$ and $|\mathrm{Im}(\gamma)/\mathrm{Re}(\gamma)|\ll 1$. For lower temperature, where higher order terms of the free energy start to matter, interstitial local minima of the free energy develop, which can be attributed to the next term in the GL expansion that features a $\phi$ dependence
$F^{(6)}=\eta(\Delta^{K}\Delta^{K}\Delta^{K}\Delta^{K}\Delta^{K}\Delta^{K}+\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}})\propto\mathrm{Re}(\eta)\cos(6\phi)-\mathrm{Im}(\eta)\sin(6\phi)\ .$ (SM 13)

### v2

@@ B Causes of Fermi surface broadening in kinetic and interacting models
Figure SM 5: Fermi surface nesting of the breathing configuration with nesting vector $\boldsymbol{q}=K$ at HOVHS filling with different broadening values $T$. When $T$ becomes infinitesimal the nesting condition vanishes.
@@
Within these non-interacting models, upon lowering the FS broadening, the $\Gamma$-peak in the particle-hole susceptibilty surpasses the apex at $\boldsymbol{q}=K$. Visualizing the FS overlap of a $K$-shifted FS for different values of temperature broadening $T$ (Fig. SM 5) in the breathing configuration elucidates this observation. In the absence of broadening, i.e. a weak-coupling scenario, the overlaps vanish because of non-matching curvatures of the FS regions.
@@
In the case of $K$-ordering, the appropriate symmetry group is no longer $C_{3v}$ of the original lattice, but the extended pointgroup $C_{3v}^{\prime\prime}$ of the enlarged $\sqrt{3}\times\sqrt{3}$ unit cell. When comparing the transformation behavior of the two states in Fig. SM 6 with the character table Table SI, the order parameter can be labeled by the $E_{3}^{\prime}$ irrep.
Figure SM 6: Degenerate leading FRG eigenvectors in the Kekulé phase at $K$ and $K^{\prime}$} up to nearest neighbor bond distances. The size (color) of the sites and bonds indicate the absolute value (complex phase) of the order parameter $\Delta^{K^{(\prime)}}_{o_{1}o_{2}}(\bm{r})=\frac{1}{N_{\bm{k}}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{r}}\,\Delta^{K^{(\prime)}}_{o_{1}o_{2}}(\bm{k})$. The original (extended) Wigner-Seitz cell is indicated in grey (black).
For a symmetry group with multiple equivalent rotation centers like $C_{3v}$ (equivalent to the wallpaper group $P3m1$), the angular momentum for the center of mass coordinate of the order becomes dependent on the choice of the rotation center. E.g. if we choose the $C_{3}$ center in Fig. SM 6 as the center of a strong bond triangle, the obtained order acquires a trivial transformation behavior under rotation in contrast to the phase winding around the center of the hexagon.
Table SI: Character table of the extended pointgroup $C_{3v}^{\prime\prime}$ of the enlarged $\sqrt{3}\times\sqrt{3}$ unit cell of the breathing kagome lattice. $E$ is the identity operation, $t_{a_{1,2}}$ label translations with the primitive lattice vectors of the original kagome lattice, $C_{3}^{(2)}$ positive (negative) rotations by $2\pi/3$ and $\sigma_{1,2,3}$ are the three symmetry equivalent mirrors through the sides of each small triangle. We use the center of the hexagonal plaquettes as the symmetry center and note that a different rotation center leads to a permutation of the characters of the irreps $E_{i}^{\prime}$.
@@
$\bar{\Delta}^{K}_{o_{1}o_{2}}(\bm{k})=\Delta^{{K^{\prime}}}_{o_{2}o_{1}}(\bm{k}+K)\ .$ (SM 5)
@@
From the FRG analysis, we know that the non-vanishing expectation value in the symmetry broken state is given by $\langle{\hat{c}^{\dagger}}_{\bm{k}+\bm{q}}{\hat{c}^{\vphantom{{\dagger}}}}_{k}\rangle$ with $\bm{q}\in\{K,{K^{\prime}}\}$ and we can write down the corresponding mean field decoupling for the interacting Hamiltonian
[…]
@@
$\Delta^{\bm{q}}_{o_{1}o_{2}}(\bm{k})=\frac{1}{N_{\bm{k}}}\sum_{{\bm{k}^{\prime}}o_{3}o_{4}}V^{D}_{o_{1}o_{2}o_{3}o_{4}}(\bm{q},\bm{k},{\bm{k}^{\prime}})\langle{\hat{c}^{\dagger}}_{{\bm{k}^{\prime}}-\bm{q}o_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{{\bm{k}^{\prime}}o_{4}}\rangle\,.$ (SM 7)
@@
$\begin{split}C={}&-\frac{1}{N_{\bm{k}}}\sum_{\bm{q}\in\{K,{K^{\prime}}\}}\sum_{\bm{k}{\bm{k}^{\prime}}\{o_{i}\}}V^{D}_{o_{1}o_{2}o_{3}o_{4}}(\bm{q},\bm{k},{\bm{k}^{\prime}})\langle{\hat{c}^{\dagger}}_{\bm{k}+\bm{q}o_{1}}{\hat{c}^{\vphantom{{\dagger}}}}_{\bm{k}o_{2}}\rangle\langle{\hat{c}^{\dagger}}_{{\bm{k}^{\prime}}-\bm{q}o_{3}}{\hat{c}^{\vphantom{{\dagger}}}}_{{\bm{k}^{\prime}}o_{4}}\rangle\\ ={}&-\frac{1}{N_{\bm{k}}}\sum_{\bm{q}\in\{K,{K^{\prime}}\}}\sum_{\bm{k}{\bm{k}^{\prime}}\{o_{i}\}}\big(V^{D}_{o_{1}o_{2}o_{3}o_{4}}(\bm{q},\bm{k},{\bm{k}^{\prime}})\big)^{-1}\Delta_{o_{1}o_{2}}^{-\bm{q}}(\bm{k})\Delta_{o_{3}o_{4}}^{\bm{q}}({\bm{k}^{\prime}})\end{split}$ (SM 8)
@@
Using the spinor notation ${\hat{\psi}^{\dagger}}_{\bm{k}o_{1}}=\big({\hat{c}^{\dagger}}_{\bm{k}o_{1}},{\hat{c}^{\dagger}}_{\bm{k}+Ko_{1}},{\hat{c}^{\dagger}}_{\bm{k}+{K^{\prime}}o_{1}}\big)$, the MF Hamiltonian can be written as
$H_{\mathrm{MF}}=\frac{1}{3}\sum_{\bm{k}o_{1}o_{2}}{\hat{\psi}^{\dagger}}_{\bm{k}o_{1}}\begin{pmatrix}\mathcal{H}^{0}_{o_{1}o_{2}}(\bm{k})&2\Delta_{o_{1}o_{2}}^{{K^{\prime}}}(\bm{k}+K)&2\Delta_{o_{1}o_{2}}^{K}(\bm{k}+{K^{\prime}})\\ 2\Delta_{o_{1}o_{2}}^{K}(\bm{k})&\mathcal{H}^{0}_{o_{1}o_{2}}(\bm{k}+K)&2\Delta_{o_{1}o_{2}}^{{K^{\prime}}}(\bm{k}+{K^{\prime}})\\ 2\Delta_{o_{1}o_{2}}^{{K^{\prime}}}(\bm{k})&2\Delta_{o_{1}o_{2}}^{K}(\bm{k}+K)&\mathcal{H}^{0}_{o_{1}o_{2}}(\bm{k}+{K^{\prime}})\end{pmatrix}{\hat{\psi}^{\vphantom{{\dagger}}}}_{\bm{k}o_{2}}+C=\frac{1}{3}\sum_{\bm{k}n}\xi_{n}(\bm{k}){\hat{\psi}^{\dagger}}_{\bm{k}n}{\hat{\psi}^{\vphantom{{\dagger}}}}_{\bm{k}n}+C\,$, (SM 9)
and we obtain the quasi-particle spectrum $\xi_{n}(\bm{k})$ within the ordered phase by diagonalizing the Hamiltonian. We note that the symmetry constraint of Eq. SM 5 guarantees a hermitian MF Hamiltonian. Exploiting the fermionic nature of the quasi-particles, i.e., ${\hat{\psi}^{\dagger}}_{\bm{k}n}{\hat{\psi}^{\vphantom{{\dagger}}}}_{\bm{k}n}=\{0,1\}$, the free energy can by directly calculated as
$F=-T\ln(Z)=-T\ln\Big(\prod_{\bm{k}n}\sum_{n_{\bm{k}n}=0}^{1}e^{-\frac{\xi_{n}(\bm{k})}{3T}}\Big)+C=-T\sum_{\bm{k}n}\ln\Big(1+e^{-\frac{\xi_{n}(\bm{k})}{3T}}\Big)+C\,.$ (SM 10)
@@
$\Delta_{o_{1}o_{2}}(\bm{k})=e^{i\phi_{K}}A_{K}\Delta^{K}_{o_{1}o_{2}}(\bm{k})+e^{i\phi_{{K^{\prime}}}}A_{{K^{\prime}}}\Delta^{{K^{\prime}}}_{o_{1}o_{2}}(\bm{k})\,.$ (SM 11)
Since Eq. SM 5 is not enforced by TRS directly, but is rather a direct consequence of hermiticity and $K$ and $K^{\prime}$} being time reversal partners, this requirement persists in the symmetry broken phase. The order parameter in the $K$-phase features a finite site-local, and hence momentum independent, component. Thereby Eq. SM 5 directly implies $A_{K}=A_{{K^{\prime}}}$ and $\phi_{K}=-\phi_{{K^{\prime}}}$ and we need to optimize the free energy with respect to the simplified order parameter
$\Delta_{o_{1}o_{2}}(\bm{k})=A\big[e^{i\phi}\Delta^{K}_{o_{1}o_{2}}(\bm{k})+e^{-i\phi}\Delta^{{K^{\prime}}}_{o_{1}o_{2}}(\bm{k})\big]\,.$ (SM 12)
A MF analysis after an FRG flow technically requires a regularization of the divergent part of the two particle vertex via an inverse Bethe-Salpeter resummation [114, 80]. With this approach, all degrees of freedom are restored and the correct symmetry breaking on the mean field level is ascribed to a modified bare interaction. As apparent from Eq. SM 6, the effective two-particle interaction only enters the term $C$, which merely determines the absolute amplitude of the order parameter as a function of temperature, but cannot distinguish between different linear combinations within the degenerate manifold. Hence, we only need to evaluate the kinetic term in Eq. SM 10 as a function of $\phi$ for different values of $T\lesssim T_{\mathrm{C}}$ and $A$. This can be done for the original theory without any further restrictions [87].
The free energy in Fig. SM 7 shows three distinct minima at $\phi\in\{\pi/3,\pi,5\pi/3\}$ for all values of $A$ and $T$, which are related by a translation by an original primitive translation. Hence, all minima can be mapped back to the same state by a redefinition of the unit cell. The obtained state is depicted in Fig. 2b of the main text.
Figure SM 7: Phase dependence of the free energy for different inverse temperatures $\beta=1/T$ at fixed amplitude $A=t$ (solid lines). $F$ was calculated according to Eq. SM 10 on a $1000\times 1000$ momentum grid in the downfolded BZ. The dependence can be fitted by taking into account third and sixth order terms in the GL expansion in Eq. SM 15(dashed lines).
@@ A Complementary Ginzburg-Landau analysis
$F=\alpha\Delta^{K}\Delta^{K^{\prime}}+\gamma(\Delta^{K}\Delta^{K}\Delta^{K}+\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}})+\mathcal{O}(\Delta^{4})\ ,$ (SM 13)
where the summation over the internal momentum structure of $\Delta^{K^{(\prime)}}(\bm{k})$ has been carried out and incorporated in the GL coefficients $\alpha-\gamma$. While $\alpha\propto(T-T_{\mathrm{C}})$ determines the absolute size of the order parameter $A$ close to $T_{\mathrm{C}}$, the third order term $\propto\mathrm{Re}(\gamma)\cos(3\phi)-\mathrm{Im}(\gamma)\sin(3\phi)$ is responsible for fixing the overall phase of the condensate [7].
For $\beta=1/T=10$, i.e. small size of the order parameter, where the expansion in Eq. SM 13 is still valid, the phase dependence of the free energy can be appropriately fitted by setting $\mathrm{Re}(\gamma)>0$ and $|\mathrm{Im}(\gamma)/\mathrm{Re}(\gamma)|\ll 1$. For lower temperature, where higher order terms of the free energy start to matter, interstitial local minima of the free energy develop, which can be attributed to the next term in the GL expansion that features a $\phi$ dependence
$F^{(6)}=\eta(\Delta^{K}\Delta^{K}\Delta^{K}\Delta^{K}\Delta^{K}\Delta^{K}+\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}}\Delta^{K^{\prime}})\propto\mathrm{Re}(\eta)\cos(6\phi)-\mathrm{Im}(\eta)\sin(6\phi)\ .$ (SM 14)